# DualSG: A Dual-Stream Explicit Semantic-Guided Multivariate Time Series Forecasting Framework

### v3

@@ 1. Introduction
While the dual streams in DualSG capture complementary aspects of time series, modeling variable-wise dependencies remains challenging. Existing methods, including full inter-channel attention (Zhang and Yan, 2023; Shi et al., 2025), channel-wise independence (Liu et al., 2024d; Nie et al., 2023), and clustering-based grouping (Qiu et al., 2025; Liu et al., 2024a), involve trade-offs among scalability, expressiveness, and adaptability, and often lack semantic reasoning. To address this, we propose SemFuse, a lightweight fusion module that leverages TSCs to guide sparse and interpretable inter-channel interactions via trend-level semantic similarity.
To integrate the numerical and textual streams, we introduce the Spatial & Temporal Attention Matrix (STAM), which dynamically adjusts their contributions across time and variables. Instead of uniform fusion, STAM performs context-aware weighting to emphasize either fine-grained signals or semantic corrections as needed, enhancing both short-term accuracy and long-term trend consistency under complex or distribution-shifted scenarios.
@@ 3. Methodology
3.1. Preliminaries
Multivariate Time Series Forecasting (MTSF) aims to predict the next $F$ future time steps, given a historical multivariate time series. Formally, let the input sequence be denoted as $\mathbf{X}=\langle\mathbf{x}_{:,1},\ldots,\mathbf{x}_{:,T}$$\rangle\in\mathbb{R}^{N\times T}$, where $N$ is the number of variables (channels) and $T$ is the length of the historical window. The forecasting objective is to estimate the future sequence $\mathbf{Y}=\langle\mathbf{x}_{:,T+1},\ldots,\mathbf{x}_{:,T+F}\rangle\in\mathbb{R}^{N\times F}$.
@@ 3.3. Textual Reasoning Stream
Time-to-Text Cross Attention. To generate natural language captions from time series, we adopt a cross-attention mechanism where decoder tokens attend to time series features as conditional context.
Formally, given the encoded time series $H_{\text{time}}\in\mathbb{R}^{T\times d}$, we apply cross attention at each decoder layer. For each decoding step, the token embeddings $H_{\text{text}}\in\mathbb{R}^{L\times d}$ serve as queries:
(11) $Q=H_{\text{text}}W_{Q},\quad K=H_{\text{time}}W_{K},\quad V=H_{\text{time}}W_{V}$
(12) $H_{\text{cross}}=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V$
(13) $H_{\text{align}}=\text{LayerNorm}(H_{\text{text}}+H_{\text{cross}})$
This enables the decoder to extract relevant temporal semantics during generation. Unlike classical encoder-decoder frameworks, our design injects time-conditioned signals directly into each layer of the GPT2 decoder, allowing flexible integration with pre-trained language models.
@@
Figure 3. Overview of the TSCG module. The model embeds time series patches with positional encoding, extracts temporal features via a Transformer encoder, aligns features with a textual template using cross attention, and generates captions through an autoregressive decoder.
@@ 3.4.1. Channel Fusion via SemFuse
Efficiency and Robustness. Compared to Channel Dependency (CD) methods that compute full pairwise attention, and Channel Clustering (CC) approaches that enforce static groupings, SemFuse is more efficient and flexible. By fusing channels through a one-shot, caption-guided pathway, SemFuse avoids exhaustive computation while retaining interpretability. Since fusion occurs at the semantic level, it is also more robust to local fluctuations in raw inputs. We summarize the computational complexity of various channel fusion methods in Table 2.
Table 2. Computational complexity comparison among different channel fusion strategies. NN denotes the number of channels, and $K$ denotes the number of selected channels for $C$ methods.
Method Computational Complexity Operation Description
CD $O(N^{2})$ Compute pairwise attention weights and weighted sum
$C$ $O(NK)$ Compute attention weights within clusters ($K<N$) and weighted sum
SemFuse (ours) $O(N)$ Concatenate caption-based selected channels directly
@@ 4.2. Main Results
Statistics Input statistics: a) Min value: $x_{\text{min}}$ b) Max value: $x_{\text{max}}$ c) Median value: $x_{\text{median}}$ d) Trend: upward or downward e) Top 5 lags: $x_{\text{top}}$
@@ 4.3.2. Comparison of Prompt Strategies
To explore the effectiveness of different textual inputs, we benchmark six prompt types under the same forecasting framework: domain, timestamp, statistics, instruction, external, and our proposed caption, as shown in Table 4. As illustrated in Figure  4, caption-based prompts consistently outperform others, achieving the lowest average MSE and the highest frequency of best results. The timestamp prompt ranks second, aligning with prior studies that highlight the benefit of temporal markers. Other prompt types suffer from lower semantic richness or contextual relevance. The external prompt category was excluded due to the absence of timestamp-aligned paired datasets.
@@ 4.3.3. Limitations of Fixed-length Patching and Different LLMs
Table 7. Full results for long-term forecasting consider prediction horizons $H$ within $\{96,192,336,512\}$. The lookback windows search from $\{96,192,336,512,720\}$. The term ‘Avg.” represents the average results across the four prediction lengths. The best and second best outcomes are highlighted in best and second, respectively. The notation $1^{\text{st}}$ Count” denotes the frequency of each method achieving the top results.
@@ Acknowledgements.
This work was supported by the Innovation Funding of ICT, CAS under Grant No. E461070.

### v4

@@ 1. Introduction
While the dual streams in DualSG capture complementary aspects of time series, modeling variable-wise dependencies remains challenging. Existing methods, including full inter-channel attention (Zhang and Yan, 2023; Shi et al., 2025), channel-wise independence (Liu et al., 2024d; Nie et al., 2023), and clustering-based grouping (Qiu et al., 2025; Liu et al., 2024a), involve trade-offs among scalability, expressiveness, and adaptability, and often lack semantic reasoning. To address this, we propose SemFuse, a lightweight fusion module that leverages TSCs to guide sparse and interpretable inter-channel interactions via trend-level semantic similarity. To integrate the numerical and textual streams, we introduce the Spatial & Temporal Attention Matrix (STAM), which dynamically adjusts their contributions across time and variables. Instead of uniform fusion, STAM performs context-aware weighting to emphasize either fine-grained signals or semantic corrections as needed, enhancing both short-term accuracy and long-term trend consistency under complex or distribution-shifted scenarios.
@@ 3.2. Textual Reasoning Stream
Time-to-Text Self Attention. To generate natural language captions directly from time series, we employ a multi-head self-attention mechanism that maps temporal representations into textual space. Different from standard encoder-decoder cross attention, our design does not rely on external textual queries; instead, the decoder operates solely on time-series features.
Formally, given the encoded time series $H_{\text{time}}\in\mathbb{R}^{T\times d}$, we compute the attention states as:
(11) $Q=H_{\text{time}}W_{Q},\quad K=H_{\text{time}}W_{K},\quad V=H_{\text{time}}W_{V}$
(12) $H_{\text{attn}}=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V$
(13) $H_{\text{align}}=\text{LayerNorm}(H_{\text{time}}+H_{\text{attn}})$
This formulation directly injects temporal signals into the GPT2 decoder, allowing it to gradually transform time-series dynamics into linguistic representations. By avoiding mixed-modal cross-attention, the model ensures that the textual output is conditioned purely on numerical patterns, while still benefiting from the flexibility of pre-trained language models.
@@ 3.3.1. Channel Fusion via SemFuse
Efficiency and Robustness. Compared to Channel Dependency (CD) methods that compute full pairwise attention, and Channel Clustering (CC) approaches that enforce static groupings, SemFuse is more efficient and flexible. By fusing channels through a one-shot, caption-guided pathway, SemFuse avoids exhaustive computation while retaining interpretability. Since fusion occurs at the semantic level, it is also more robust to local fluctuations in raw inputs.
@@ 4.2. Main Results
Statistics Input statistics: a) Min value: $x_{\text{min}}$ ; b) Max value: $x_{\text{max}}$; c) Median value: $x_{\text{median}}$; d) Trend: upward or downward; e) Top 5 lags: $x_{\text{top}}$
@@ 4.3.2. Comparison of Prompt Strategies
To explore the effectiveness of different textual inputs, we benchmark six prompt types under the same forecasting framework: domain, timestamp, statistics, instruction, external, and our proposed caption, as shown in Table 3. As illustrated in Figure  3, caption-based prompts consistently outperform others, achieving the lowest average MSE and the highest frequency of best results. The timestamp prompt ranks second, aligning with prior studies that highlight the benefit of temporal markers. Other prompt types suffer from lower semantic richness or contextual relevance. The external prompt category was excluded due to the absence of timestamp-aligned paired datasets.
@@ 4.3.3. Limitations of Fixed-length Patching and Different LLMs
Table 6. Full results for long-term forecasting consider prediction horizons $H$ within $\{96,192,336,512\}$. The lookback window for our framework is selected via a search over $\{96,192,336,512,720\}$. The best and second best outcomes are highlighted in best and second, respectively. The notation $1^{\text{st}}$ Count” denotes the frequency of each method achieving the top results.
@@ Acknowledgements.
This work was supported by the Innovation Funding of Institute of Computing Technology, Chinese Academy of Sciences under Grant No. E461070.